# On How Iterative Magnitude Pruning Discovers Local Receptive Fields in Fully Connected Neural Networks

### v1

@@ 3.2 IMP increases non-Gaussian statistics of FCN representations
Measuring non-Gaussian statistics in neural networks representations. Because the non-Gaussian statistics, present in natural images, contain information that is localized, a neural network that develops local RFs should increasingly represent these non-Gaussian statistics [24]. To quantitatively probe this, we can measure the “non-Gaussanity” of a given neural network’s representations. One such way of doing this is by computing the statistical kurtosis of the preactivations. We chose to use the kurtosis (as opposed to another metric), as it was previously shown by Ingrosso and Goldt (2022) [24] that the kurtosis significantly increased with training as FCNs developed localized RFs.
The activation of units in the first hidden layer of the neural network, $f(\theta;X)$, for an image $X\in\mathbb{R}^{N_{p}^{2}}$, are given by
$a_{i}=\sigma\left(\sum_{j=1}^{N_{p}^{2}}W_{ij}X_{j}+b_{i}\right),$ (2)
@@
$\text{kurt}(\lambda_{i})=\frac{\mathbb{E}_{X}\left[\lambda_{i}(X)-\mathbb{E}_{X}\lambda_{i}(X)\right]^{4}}{\left(\left[\lambda_{i}(X)-\mathbb{E}_{X}\lambda_{i}(X)\right]^{2}\right)^{2}},$ (4)
@@
where $\mathbb{E}_{X}$ is the expectation computed over all the inputs $X\in\mathcal{X}$. This can be practically computed by considering all $X$ in the training or test sets. The kurtosis allows us to quantify how non-Gaussian the distribution of preactivations is. If $\lambda_{i}$ is distributed according to a Gaussian, then $\text{kurt}(\lambda_{i})=3$. If $\text{kurt}(\lambda_{i})>3$, then the distribution is more peaked than a Gaussian. If $\text{kurt}[\lambda_{i}]<3$, the distribution is more broad than a Gaussian. The extent to which a distribution of preactivations is non-Gaussian can measured by its excess kurtosis, defined as $\left|3-\text{kurt}(\lambda_{i})\right|$.
@@
IMP increases FCN preactivation kurtosis, relative to oneshot pruning. Given that IMP leads to more localized RFs than oneshot pruning (Fig. 1) [17], we hypothesize that it increases the non-Gaussian statistics being represented by the FCN more strongly than oneshot pruning. In particular, because IMP is applied iteratively, we hypothesize that each round of pruning leads its starting point, $f[\theta(t_{\text{rewind}})\odot m(n-1);X]$, to have a greater representation of non-Gaussian statistics, providing a kind of feedback loop that drives the emergence of more localized RFs.
@@
To test this, we compute the kurtosis of the preactivations of $f[\theta(t_{\text{rewind}})\odot m(n-1);X]$, for the masks found by both IMP and oneshot pruning. We also compare with a random pruning baseline, to identify how much the observed change in kurtosis is due sparsity alone. As expected, we find that both IMP and oneshot pruning discover masks with significantly more localized structure than random pruning (Fig. 4A). In-line with non-Gaussian statistics driving localization, we find significant increases in kurtosis only for IMP and oneshot pruning (Fig. 4B). This demonstrates that the increase in local non-Gaussian statistics is not merely a product of sparsification, as random pruning does not see this effect.
@@
Figure 4: IMP increases preactivation kurtosis and localization of RFs in FCNs trained on ImageNet32, to a greater extent than oneshot pruning. (A) Mean RF width, as a function of sparsity induced by either IMP (black line), oneshot pruning (blue line), or random pruning (red line). (B) Mean kurtosis of preactivations, per class, as a function of sparsity induced by IMP, oneshot pruning, and random pruning. Note that a kurtosis $>3$ implies more non-Gaussian statistics. In (A)-(B), solid line is mean and shaded area is minimum and maximum of three independently trained and pruned FCNs.
@@
When comparing IMP and oneshot pruning, we find that IMP leads to greater preactivation kurtosis (Fig. 4B), at a sparsity that precedes IMP’s greater localization (Fig. 4A). In addition, we find that the preactivation kurtosis increases almost monotonically for IMP. These results support our hypothesis that IMP increases the non-Gaussanity of its representations with each round of pruning, to a greater extent than would be achieved by oneshot pruning. Thus, the iterative nature of IMP is playing an important role in amplifying the non-Gaussian statistics.
@@ 3.3 IMP maximally increases non-Gaussian statistics of FCN representations
Probing the role of individual weights on the statistics of FCN representations. While the results presented in Fig. 4 demonstrate that the sparse subnetworks IMP finds after each round of pruning, $f[\theta(t_{\text{rewind}})\odot m(n-1);X]$, start with increasingly non-Gaussian representations, as compared to oneshot pruning, they do not provide insight on whether the parameters IMP removes are “optimal” in driving the largest increase in non-Gaussanity. Evidence for this would directly support our central hypothesis (H).
Because IMP removes multiple parameters at once, identifying the choice that maximizes the non-Gaussian representation is combinatorially explosive, and thus, is not tractable. We can none-the-less approach testing this hypothesis by considering a simplified setting. In particular, we take inspiration from the “Cavity Method” in statistical physics [52], and evaluate the impact a given weight in the first hidden layer, $W_{ij}$, has on the statistics of the preactivations, $\lambda_{i}$, by computing the kurtosis of the preactivations, with and without the weight [^2] We consider only weights here, as removing the biases does not change the kurtosis of the preactivations, as it is just corresponds to a translation of the distribution.. To quantify weight $W_{ij}$’s impact on the statistics, we develop a metric we call a “cavity score”, computed as
@@
A negative cavity score signifies that removing $W_{ij}$ decreases the non-Gaussanity of the preactivations (making them more Gaussian), and a positive cavity score signifies that removing $W_{ij}$ increases the non-Gaussanity of the preactivation (making them less Gaussian). A schematic of the approach is shown in Fig. 5A.
@@
Figure 5: IMP selectively prunes weights when their removal would most increase the non-Gaussianity of the preactivations. (A) A schematic overview of how the cavity score is computed. For a given unit in the first hidden layer, the kurtosis of its preactivation is computed (top). Then, the preactivation kurtosis is recomputed, with a given weight $W_{ij}$ removed (bottom). If the distribution of preactivations becomes more Gaussian once $W_{ij}$ is removed, $\text{cavity}(W_{ij})<0$. If the distribution of preactivations instead becomes less Gaussian, $\text{cavity}(W_{ij})>0$. (B) Mean cavity score, computed at IMP round 0, for weights grouped according to the round of IMP they ultimately get removed during. Note that the weights that get removed later during IMP have negative cavity score, while weights that get removed early during IMP have positive cavity score. (C) Same as (B), but when computing the cavity score of the remaining weights, $\theta(t_{\text{rewind}})\odot m(n-1)$, after each round of IMP. Gray dashed line is used to highlight the fact that the mean cavity score of the weights that get removed at IMP round 8 is negative for all rounds of IMP, until after ${}^{\text{th}}$ round of pruning. In (B)-(C), solid line denote mean, and shaded area is minimum and maximum of three independently trained and pruned FCNs.
@@
The order in which IMP prunes maximizes the representation of non-Gaussian statistics in FCNs. We use the cavity score to probe whether the ordering with which IMP removes weights selectively shapes the statistics of the preactivation kurtosis to be less Gaussian. Re-analyzing the IMP experiments used in Fig. 4, we compute the cavity score for all weights, based on their value at the rewind point [$\theta(t_{\text{rewind}})$ – before any pruning has been performed, i.e. IMP round 0]. We then group the weights by which round of IMP they ultimately get pruned during [^3] The re-analysis allows us to do this, as we have already run the experiment, and therefore know the “fate” of each weight.. Plotting the average cavity score, for each grouping, we see that the weights that get removed in the first and second round of IMP have positive cavity score, while all weights that get removed during IMP round 4 or later have negative cavity score (with the weights that get removed at IMP round 3 having cavity score approximately equal to zero) (Fig. 5B). Thus, for a weight that gets removed during the later rounds of IMP ($n\geq 4$), its removal at this initial stage would decrease in the non-Gaussianity of the preactivations. In contrast, for a weight that is removed during the earliest rounds of IMP, its removal at this initial stage would increase the non-Gaussanity of the preactivations.
@@
We then consider whether this same trend holds for the subsequent rounds of IMP. As the preactviation distribution for each hidden unit changes after pruning (as some weights get removed), we recompute the cavity score for all the remaining weights after the first round of IMP (at the start of IMP round $n=2$). As before, the preactivations are measured from the rewind point [now with the parameters $\theta(t_{\text{rewind}})\odot m(1)$]. We again group the remaining weights by the round of IMP they get removed during, and plot the mean. This process is repeated for all subsequent rounds of IMP. Because the number of weights remaining after each round of IMP decreases, the removal of any given surviving weight leads to an increasingly large impact on the preactivations. To avoid seeing trivial differences in scaling, we report $N_{W}(n)\cdot\text{cavity}(W_{ij})$, where $N_{W}(n)=N(1-s)^{n}$ is the number of weights remaining after $n$ rounds of IMP. This acts as a form of normalization to enable more fair comparison, but it does not affect our primary conclusion, which rests on the sign of the cavity score.
@@
That we see this trend of for all rounds of IMP strongly supports the idea that it is the order, in addition to the identity, of the weights that IMP removes, that leads to an increase in the non-Gaussian preactivation statistics. Further, this order maximizes the amount of non-Gaussanity of the FCN representation, as IMP does not remove weights when their pruning would increase the Gaussian statistics of the preactivations. This provides strong support for our central hypothesis, H.
@@ 4 Discussion
Motivated by finding that iterative magnitude pruning discovers local receptive fields in fully connected neural networks [17], we sought to investigate what the cause of this phenomenon is. By connecting Pelligrini and Biroli’s (2022) [17] work to with Ingrosso and Goldt’s (2022) [24] demonstration that local RFs can emerge in FCNs trained on inputs with sufficiently strong non-Gaussian statistics, we hypothesized that IMP iteratively maximizes the the amount of non-Gaussian statistics present in the FCN preactivation, at each round of pruning (Hypothesis H).
To support this hypothesis, we made three observations. First, we showed that non-Gaussian statistics are necessary for local RFs to emerge via IMP, as training FCNs on datasets that match ImageNet32 in the first two cumulants, and have zero higher-order cumulants, leads to IMP finding masks with no localization (Fig. 3). Second, we found that IMP amplifies the amount of non-Gaussanity present in the FCN representation with each round of pruning, making the distribution of preactivations significantly more non-Gaussian than oneshot pruning (Fig. 4). And third, by developing a “cavity method” inspired approach, we provided evidence that the order in which IMP removes FCN weights is aligned with when their removal would increase the non-Gaussanity of the preactivations (Fig. 5). Collectively, these provide strong evidence for H.
@@
This work is focused on shallow FCNs, trained on a down-sampled version of ImageNet (ImageNet32 [49]). While this is a setting that is removed from many modern implementations of machine learning (ML), we believe this is justified by the fact that it is the only setting in which the inductive bias discovered by IMP is well characterized [17]. To fully prove our central hypothesis (H), it is necessary to show that the weights IMP removes are optimal in increasing the preactivation kurtosis, over all possible choices of pruning. This is infeasible to show. However, our cavity method (Sec. 3.3) is able to provide evidence of a first order approximation of optimality of IMP (Fig. 5). We believe this, along with our results, provide strong evidence for the validity of H.
@@
Recent work examining the training dynamics that drive local RF emergence in a simplified batch gradient setting, with synthetic data and a single hidden layer FCN, argue that training on inputs with kurtosis $<3$ (negative excess kurtosis) is essential to their formation [25]. This is in contrast to our results on FCNs trained on ImageNet32, where we found localization in pruning masks, despite the preactivation kurtosis being $>3$ (Fig. 1). This difference may point to implicit properties of stochastic gradient descent and its interaction with large-scale computer vision datasets, which future work can explore. That the order in cavity score is well aligned to the order of when weights are removed via IMP, before any pruning occurs (Fig. 5C – IMP round 0 line), suggests that it could be used to more efficiently find localized RFs. Future work can explore how to best use this metric as a cost function to develop new pruning algorithms.
Lessons for sparse ML.
In addition to IMP, other approaches have been developed to identify sparse subnetworks. These include methods that oneshot prune dense models based on gradients and Hessians of the loss [53, 54, 55], and dynamic sparsity methods that optimize over masks [56, 57], thereby never training on the dense model. Our results suggest that the ability of IMP to identify local RFs in FCNs comes both from its iterative nature and its leveraging of the dense model’s training dynamics. This calls into question whether these other sparsification approaches are capable of identifying local RFs, and has implications on the ability of these approaches to extract and maintain strong inductive biases.
@@
While FCNs offer a tractable framework with which to study the impact of IMP on internal representations, the success of IMP in modern architectures, such as CNNs [6, 3, 12], suggests that IMP may similarly be increasing the non-Gaussian statistics of the preactivations and driving the emergence of even stronger local RFs. Future work can explore the effect of IMP on the non-Gaussian statistics in CNNs. IMP has also been successfully applied in settings beyond computer vision classification tasks, including natural language processing [36, 11] and reinforcement learning [36, 40]. In such settings, useful inductive biases and their interaction with the statistical properties of the inputs are less clear. An exciting avenue of future work would be to leverage the cavity method developed in this work to explore the impact of individual weights and the order with which IMP removes them. Work in this direction may shed additional light on why IMP can struggle to sparsify large language models [58, 59].
Appendix A PCA and ICA Experiments
We compute principal component analysis (PCA) and independent component analysis (ICA) [48] using

### v3

@@ 3.2 IMP increases non-Gaussian statistics of FCN representations
Because the non-Gaussian statistics present in natural images contain information that is localized, a neural network that develops local RFs should increasingly represent these non-Gaussian statistics [26]. To quantitatively probe this, we can measure the “non-Gaussanity” of a given neural network’s representations. One such way of doing this is by computing the excess kurtosis of the pre-activations, as we describe below.
The pre-activations of fully-connected neural networks
The activation of units in the first hidden layer of the neural network, $f(\theta;\mathcal{X})$, for an image $X\in\mathcal{X}$, are given by
$a_{i}(X)=\sigma\left(\sum_{j=1}^{N_{p}^{2}}W_{ij}X_{j}+b_{i}\right),$ (2)
@@
Measuring non-Gaussian statistics in neural networks representations.
@@
$\text{kurt}(\lambda_{i})=\frac{\mathbb{E}_{\mathcal{X}}\left[\lambda_{i}(X)-\mathbb{E}_{\mathcal{X}}\lambda_{i}(X)\right]^{4}}{\left(\left[\lambda_{i}(X)-\mathbb{E}_{\mathcal{X}}\lambda_{i}(X)\right]^{2}\right)^{2}},$ (4)
@@
where $\mathbb{E}_{X}$ is the expectation computed over all the inputs $X\in\mathcal{X}$. The kurtosis allows us to quantify how non-Gaussian the distribution of preactivations is. If $\lambda_{i}$ is distributed according to a Gaussian, then $\text{kurt}(\lambda_{i})=3$. If $\text{kurt}(\lambda_{i})>3$, then the distribution is more peaked than a Gaussian. If $\text{kurt}[\lambda_{i}]<3$, the distribution is more broad than a Gaussian. The extent to which a distribution of preactivations is non-Gaussian can measured by its excess kurtosis, defined as $\left|3-\text{kurt}(\lambda_{i})\right|$. We chose to use the kurtosis (as opposed to another metric), as it was previously shown by Ingrosso and Goldt (2022) [26] that the kurtosis significantly increased with training as FCNs developed localized RFs.
IMP increases FCN preactivation kurtosis, relative to oneshot pruning.
@@
Given that IMP leads to more localized RFs than oneshot pruning (Fig. 1) [17], we hypothesize that IMP increases the non-Gaussian statistics being represented by the FCN more strongly than oneshot pruning. In particular, because IMP is applied iteratively, we hypothesize that each round of pruning leads its starting point, $f[\theta(t_{\text{rewind}})\odot m(n-1);\mathcal{X}]$, to have a greater representation of non-Gaussian statistics, providing a kind of feedback loop that drives the emergence of more localized RFs.
@@
To test this, we compute the kurtosis of the preactivations of $f[\theta(t_{\text{rewind}})\odot m(n-1);\mathcal{X}]$, for the masks found by both IMP and oneshot pruning. We also compare with a random pruning baseline, to identify how much the observed change in kurtosis is due sparsity alone. As expected, we find that both IMP and oneshot pruning discover masks with significantly more localized structure than random pruning (Fig. 4A). In-line with non-Gaussian statistics driving localization, we find significant increases in the kurtosis of the preactivations only for IMP and oneshot pruning (Fig. 4B). This demonstrates that the increase in local non-Gaussian statistics is not merely a product of sparsification, as random pruning does not see this effect.
@@
Figure 4: IMP increases localization of RFs and preactivation kurtosis in FCNs trained on ImageNet32, to a greater extent than oneshot pruning. (A) Mean RF width, as a function of sparsity induced by IMP (black line), oneshot pruning (blue line), or random pruning (red line). (B) Mean kurtosis of preactivations, per class, as a function of sparsity induced by IMP, oneshot pruning, and random pruning. Note that a kurtosis $>3$ implies more non-Gaussian statistics. In (A)-(B), solid line is mean and shaded area is minimum and maximum of three independently trained and pruned FCNs.
@@
When comparing IMP and oneshot pruning, we find that IMP leads to greater preactivation kurtosis (Fig. 4B), at a sparsity that precedes IMP’s greater localization (Fig. 4A). In addition, we find that the preactivation kurtosis increases almost monotonically for IMP. This supports our hypothesis that IMP increases the non-Gaussanity of its representations with each round of pruning, to a greater extent than would be achieved by oneshot pruning. Thus, the iterative nature of IMP is playing an important role in amplifying the non-Gaussian statistics.
Pellegrini and Biroli [17] also found localization of the masks in the later layers of their FCN. The interpretation of this localization is more nuanced, as the hidden units in the later layers are not directly connected to the inputs. However, if non-Gaussianity is a key to localization, then we expect the preactivations in all layers that exhibit localization to become more non-Gaussian across rounds of IMP. Indeed, we find that the preactivation kurtosis in the second layer of our FCN also increases with IMP round (Fig. S3), demonstrating that IMP is also able to increase the non-Gaussianity in later FCN layers.
@@ 3.3 IMP maximally increases non-Gaussian statistics of FCN representations
Probing the role of individual weights on the statistics of FCN representations. While the results presented in Fig. 4 demonstrate that the sparse subnetworks IMP finds after each round of pruning, $f[\theta(t_{\text{rewind}})\odot m(n-1);\mathcal{X}]$, start with increasingly non-Gaussian representations, as compared to oneshot pruning, they do not provide insight on whether the parameters IMP removes are “optimal” in driving the largest increase in non-Gaussanity. Evidence for this would directly support our central hypothesis (H).
Because IMP removes multiple parameters at once, identifying the choice that maximizes the non-Gaussian representation is intractable due to its combinatorial complexity. We can nonetheless approach testing this hypothesis by considering a simplified setting. In particular, we take inspiration from the “cavity method” of statistical physics [55], and evaluate the impact a given weight in the first hidden layer, $W_{ij}$, has on the statistics of the preactivations $\lambda_{i}$ by computing the kurtosis of the preactivations, with and without the weight [^1] We consider only weights here, as removing the biases does not change the kurtosis of the preactivations, as it is just corresponds to a translation of the distribution.. To quantify weight $W_{ij}$’s impact on the statistics, we develop a metric we call the “cavity score”, computed as
@@
A negative cavity score signifies that removing $W_{ij}$ decreases the excess kurtosis of the preactivations (making them more Gaussian), and a positive cavity score signifies that removing $W_{ij}$ increases the excess kurtosis of the preactivations (making them less Gaussian). A schematic of the approach is shown in Fig. 5A.
@@
Figure 5: IMP selectively prunes weights when their removal would most increase the non-Gaussianity of the preactivations. (A) A schematic overview of how the cavity score (Eq. 5) is computed. For a given unit in the first hidden layer, the kurtosis of its preactivation is computed (top). Then, the preactivation kurtosis is recomputed, with a given weight $W_{ij}$ removed (bottom). If the distribution of preactivations becomes more Gaussian once $W_{ij}$ is removed, $\text{cavity}(W_{ij})<0$. If the distribution of preactivations instead becomes less Gaussian, $\text{cavity}(W_{ij})>0$. (B) Mean cavity score, computed at IMP round 0, for weights grouped according to the round of IMP they ultimately get removed during. Note that the weights that get removed later during IMP have negative cavity score, while weights that get removed early during IMP have positive cavity score. (C) Same as (B), but when computing the cavity score of the remaining weights, $\theta(t_{\text{rewind}})\odot m(n-1)$, after each round of IMP. Gray dashed line is used to highlight the fact that the mean cavity score of the weights that get removed at IMP round 8 is negative for all rounds of IMP, until after ${}^{\text{th}}$ round of pruning. In (B)-(C), solid line denote mean, and shaded area is minimum and maximum of three independently trained and pruned FCNs.
To get the most accurate estimate of the cavity score, we compute Eq. 5 using the entire ImageNet32 test set. However, this is computationally expensive. To reduce this cost, we could compute the cavity score over a subset of the test set. In particular, we could reduce the number of classes used to compute the cavity score (i.e. consider the average kurtosis across 10 classes of images, instead of 1000). This would accelerate the computation, as currently our method has to recompute the cavity score for each class separately. Future work could explore the trade-off between accuracy of computed cavity score and the number of test data points or classes used. We imagine that a large reduction can be achieved, while still maintaining accurate cavity score computations.
@@
The order in which IMP prunes maximizes the representation of non-Gaussian statistics in FCNs. We use the cavity score to probe whether the ordering with which IMP removes weights selectively shapes the preactivation excess kurtosis to be less Gaussian. Re-analyzing the IMP experiments used in Fig. 4, we compute the cavity score for all weights, based on their value at the rewind point, $\theta(t_{\text{rewind}})$ (before any pruning has been performed, i.e. IMP round 0). We then group the weights by which round of IMP they ultimately get pruned during [^2] The re-analysis allows us to do this, as we have already run the experiment, and therefore we know the “fate” of each weight.. Plotting the average cavity score, for each grouping, we see that the weights that get removed in the first and second round of IMP have positive cavity score, while the weights that get removed during IMP round 4 or later have negative cavity score (with the weights that get removed at IMP round 3 having cavity score approximately equal to zero) (Fig. 5B). Thus, for a weight that gets removed during the later rounds of IMP ($n\geq 4$), its removal at this initial stage would decrease the non-Gaussianity of the preactivations. In contrast, for a weight that is removed during the earliest rounds of IMP, its removal at this initial stage would increase the non-Gaussanity of the preactivations. We do not find that parameter sign or magnitude is strongly connected to cavity score (Fig. S4).
@@
We then consider whether this same trend holds for the subsequent rounds of IMP. Given that the preactviation distribution for each hidden unit changes after pruning (as some weights get removed), we recompute the cavity score for all the remaining weights after the first round of IMP (at the start of IMP round $n=2$). As before, the preactivations are measured from the rewind point (now determined by $f[\theta(t_{\text{rewind}})\odot m(1);\mathcal{X}]$). We again group the remaining weights by the round of IMP they get removed during, and plot the mean. This process is repeated for all subsequent rounds of IMP. Because the number of weights remaining after each round of IMP decreases, the removal of any given surviving weight leads to an increasingly large impact on the preactivations. To avoid seeing trivial differences in scaling, we report $N_{W}(n)\cdot\text{cavity}(W_{ij})$, where $N_{W}(n)=N(1-s)^{n}$ is the number of weights remaining after $n$ rounds of IMP. This acts as a form of normalization to enable more clear visualization, but it does not affect our primary conclusion, which rests on the sign of the cavity score.
@@
That we see this trend for all rounds of IMP strongly supports the idea that it is the order, in addition to the identity, of the weights that IMP removes, that leads to an increase in the non-Gaussian preactivation statistics. Furthermore, the order maximizes the amount of non-Gaussanity of the FCN representation, as IMP does not, on average, remove weights when their pruning would increase the Gaussian statistics of the preactivations. This provides strong support for our central hypothesis, H.
@@ 4 Discussion
Motivated by the finding that iterative magnitude pruning discovers local receptive fields in fully connected neural networks [17], we sought to investigate the cause of this phenomenon. By connecting this observation of Pellegrini and Biroli [17] with recent work by Ingrosso and Goldt [26], who demonstrated that local RFs can emerge in FCNs trained on inputs with sufficiently strong non-Gaussian statistics, we hypothesized that IMP iteratively maximizes the the amount of non-Gaussian statistics present in the FCN preactivation, at each round of pruning (Hypothesis H). While such a greedy algorithm may not, in general, be optimal, by creating a feedback loop (where increased non-Gaussianity leads to greater localization, which leads to greater non-Gaussianity), such a strategy could lead to strong localization.
To support this hypothesis, we made three observations. First, we showed that non-Gaussian statistics are indeed necessary for local RFs to emerge via IMP, as training FCNs on datasets that match ImageNet32 in the first two cumulants, and have zero higher-order cumulants, leads to IMP finding masks with no localization (Fig. 3). Second, we found that IMP amplifies the amount of non-Gaussanity present in the FCN representation with each round of pruning, making the distribution of preactivations significantly more non-Gaussian than oneshot pruning (Fig. 4). Third and finally, we develop a method to measure the impact of removing individual weights, inspired by the “cavity method” of statistical physics, to provide evidence that the order in which IMP removes FCN weights is aligned with when their removal would increase the non-Gaussanity of the preactivations (Fig. 5). Collectively, these results provide strong evidence for H.
@@
This work is focused on shallow FCNs, trained on a down-sampled version of ImageNet (ImageNet32 [53]). While this setting is removed from many modern implementations of machine learning (ML), we believe this is justified by the fact that it is the only setting in which the inductive bias discovered by IMP is well characterized [17]. To fully prove our central hypothesis (H), it is necessary to show that the weights IMP removes are optimal in increasing the preactivation kurtosis, over all possible choices of pruning. This is infeasible. However, our cavity method (Sec. 3.3) is able to provide evidence of a first order approximation of optimality of IMP (Fig. 5). We believe this, along with the rest of our results, provide evidence for the validity of H.
@@
Recent work examining the training dynamics that drive local RF emergence in a simplified batch gradient setting, with synthetic data and a single hidden layer FCN, argued that training on inputs with kurtosis $<3$ (negative excess kurtosis) is essential for localization [27]. This is in contrast to our results on FCNs trained on ImageNet32, where we found localization in pruning masks (Fig. 1), despite the preactivation kurtosis being $>3$ (Fig. 4B). This difference may point to implicit properties of stochastic gradient descent and its interaction with large-scale computer vision datasets, which future work can explore. That the order in cavity score is well aligned to the order of when weights are removed via IMP, before any pruning occurs (Fig. 5C – IMP round 0 line), suggests that it could be used to more efficiently find localized RFs. Future work can explore how to best use this metric as a cost function to develop new pruning algorithms.
Implications for sparse ML.
In addition to IMP, other approaches have been developed to identify sparse subnetworks. These approaches include methods that oneshot prune dense models based on gradients and Hessians of the loss [56, 57, 58], and dynamic sparsity methods that optimize over masks [59, 60], thereby never training on the dense model. Our results suggest that the ability of IMP to identify local RFs in FCNs comes both from its iterative nature and its leveraging of the dense model’s training dynamics. This may explain why Pellegrini and Biroli [17] found that SNIP [57] and SynFlow [58] were unable to find as localized RFs as IMP. Additionally, this calls into question whether dynamic sparsity training approaches are capable of identifying local RFs. More generally, our results have implications on the ability of these other sparsity algorithms to extract and maintain strong inductive biases.
@@
While FCNs offer a tractable framework with which to study the impact of IMP on internal representations, the success of IMP on modern architectures, such as CNNs [6, 3, 12], suggests that IMP may similarly be increasing the non-Gaussian statistics of the preactivations and driving the emergence of even stronger local RFs. Future work can explore the effect of IMP on the non-Gaussian statistics in CNNs. IMP has also been successfully applied in settings beyond computer vision classification tasks, including natural language processing [38, 11] and reinforcement learning [38, 42]. In such settings, useful inductive biases and their interaction with the statistical properties of the inputs are less clear. An exciting avenue of future work would be to leverage the cavity method developed in this work to explore the impact of individual weights and the order with which IMP removes them. Work in this direction may shed additional light on why IMP can struggle to sparsify large language models [61, 62].
Finally, our results also provide a possible explanation for the ability of IMP derived subnetworks (“winning tickets”) to transfer across computer vision tasks [6, 7, 8, 9, 10, 11, 12, 13] and CNN architectures [14]. In particular, if IMP increases the preactivation kurtosis in CNNs, then winning tickets will have amplified non-Gaussian statistics, which are broadly useful in computer vision tasks.
Acknowledgements
SG gratefully acknowledges funding from the European Research Council (ERC) for the project “beyond2”, ID 101166056; from the European Union–NextGenerationEU, in the framework of the PRIN Project SELF-MADE (code 2022E3WYTY – CUP G53D23000780001), and from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research” (CUP G53C22000440006).
Appendix A Measuring localization
A method for quantifying the localization of the RFs that emerges from IMP and oneshot pruning has not been developed. As a simple approximation, we choose to fit the correlation function, $S_{i}(d,n)$ (Eq. 1), with a two-dimensional Gaussian. This was motivated by the fact that Pellegrini and Biroli [17] found that the correlation function, averaged across all the most connected IMP masks, had Gaussian-like structure. As illustrated in Fig. S1A, the correlation function associated with individual IMP masks can also be well approximated by two-dimensional Gaussians. This is true even when the underlying mask does not appear to be well described by a Gaussian, thus emphasizing the rationale for fitting using $S_{i}(d,n)$. In addition, the mean-squared error (MSE) associated with the Gaussian fit of the IMP masks was significantly smaller than the MSE associated with the Gaussian fit of random pruning masks (Fig. S1B). We note that, while we believe this is evidence for the appropriateness of our method to quantify localization (via the standard deviation of the fit Gaussian), our larger point of IMP selectively increasing preactivation kurtosis is not dependent on these results.
Figure S1: Measuring localization of RFs with two-dimensional Gaussians. (A) Three example IMP masks (after 10 rounds of IMP) (left column), with their associated correlation function $S_{i}(d,n)$ (middle column). The Gaussian fits (right column) demonstrate strong qualitative agreement with the true $S_{i}(d,n)$. (B) Distribution of mean-squared error (MSE) associated with the Gaussian fits of IMP and random pruning masks. The random pruning masks serve as a baseline, as we expect them to not have much Gaussian structure in their correlation functions.
Appendix B PCA and ICA Experiments